In 1897, the first Green Jacket Golf Championship was held. The winner’s prize money was $215. In 2018, the winner’s check was $2,550,000.


a.What was the annual percentage increase in the winner’s check over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b.

If the winner’s prize increases at the same rate, what will it be in 2053? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 1,234,567.89.)
Answer #1

a. FV = PV(1 + r)t

r = (FV / PV)1/t – 1

r = ($2,550,000 / $215)1/121 – 1

r = 0.0806, or 8.06%

b. FV = PV(1 + r)t

FV = $2,550,000(1.0806)35

FV = $38,459,121.17